# A Tie-breaking based Local Search Algorithm for Stable Matching Problems

### v1

@@ 7       $M_{best}\mathrel{{:}{=}}M$;
9             $S,Q_{a}\mathrel{{:}{=}}$ Refine_Tie_Breaking_Strategy($I,M,S$);
10             $M\mathrel{{:}{=}}$ Obtain_Stable_Matching($I,M,S,Q_{a}$);
@@ 4.1 Evaluation function
where $bigM$ is a large enough number to be fixed later, $|L_{v}|$ represents the length of agent $v$’s preference list, and $b(v)-|M(v)|$ indicates the number of unassigned positions of agent $v$.
@@
where $e_{M}$ denotes the estimated minimum size of the matching throughout the entire search process. To simplify parameter settings, we define $e_{M}=c\cdot\text{size}(M_{init})$, where $c$ represents a ratio varying between 0 and 1, and $M_{init}$ denotes the matching obtained from the random tie-breaking strategy at the beginning.
@@ 4.2 Refining the tie-breaking strategy
A tie-breaking strategy, denoted as $S$, converts the original preference lists $L$, which contain ties, into tie-free (i.e., strictly ordered) preference lists $L^{*}$. Let $r^{*}_{v}(x)$ represent the rank of $x$ in $v$’s tie-free preference list. In the subroutine Refine_Tie_Breaking_Strategy, outlined in Algorithm 2, the tie-breaking strategy is updated as follows:
@@
    Obtain a set of candidate adjustments $R^{b}_{M}$ via the subroutine Obtain_Adjustments (line 3).
@@
Additionally, a set $Q_{a}$ is maintained to record the agents whose preference lists have been altered during the refinement process. This set will subsequently be used in the BP removal process within the subroutine Obtain_Stable_Matching, which is discussed in Section 4.3.
@@
Input: - An instance $I$. - A matching $M$. - A tie-breaking strategy $S$. - Information on the ties in the instance $\phi$. - A small probability $p_{d}$. - Two integer $k_{u}$ and $k_{w}$.
@@
3 Function Refine_Tie_Breaking_Strategy($I,M,S$):
@@
4       $R^{b}_{M}\mathrel{{:}{=}}$ Obtain_Adjustments($I,M$);
5       if (a small probability of $(\text{a small probability of }p_{d}\textbf{ or }R^{b}_{M}=\emptyset)$ then
@@ 8                   $Q_{a}\mathrel{{:}{=}}Q_{a}\cup v$;
[…]
10                         Break the $tie$ arbitrarily and update $S$;
@@ 18             $Q_{a}\mathrel{{:}{=}}Q_{a}\cup f$;
Next, we detail the process of identifying adjustments that can increase the likelihood of matching free agents and the construction of $R^{b}_{M}$. In the subroutine Obtain_Adjustments, as illustrated in Algorithm 3, $R^{b}_{M}$ is generated through the following two steps:
@@
    Identify a set $R_{M}$ of all potential adjustments (lines 4-11). Given a stable matching $M$, consider each free agent $f\in F_{M}$. For every agent $x\in\xi_{f}$, if the set intersection $\phi_{f}(x)\cap M(x)$ is non-empty, an adjustment is initiated. This adjustment aims to improve $f$’s rank in $x$’s tie-free preference list by positioning $f$ at the highest rank among those agents in $\phi_{f}(x)$, thereby ensuring $f$ ranks above at least one agent currently matched with $x$. This adjustment introduces $(f,x)$ as a BP and enhances $f$’s likelihood of being matched with $x$. All adjustments in $M$ are stored in $R_{M}$, which can be represented as $R_{M}=\{(f,x)\mid f\in F_{M},x\in\xi_{f},\phi_{f}(x)\cap M(x)\neq\varnothing\}$. The set of adjustments related to one free agent $f$ is denoted by $R_{M}(f)=\{a\in R_{M}\mid a[1]=f\}$, where $a[i]$ is the $i$-th element in adjustment $a$.
@@
    Construct $R^{b}_{M}$ by sampling from $R_{M}$ (lines 12-16). One free agent may have multiple adjustments, and the number of adjustments can vary among different free agents. In some cases, agents with many unassigned positions might have relatively few adjustments, while those with fewer unassigned positions might have many adjustments. This imbalance can impact the performance of the algorithm, particularly when it is difficult to increase the overall matching size by filling the positions of those agents with fewer unassigned positions. Consequently, a balanced set of adjustments, $R^{b}_{M}$, is constructed through sampling from $R_{M}$. Specifically, for each free agent $f\in F_{M}$, we randomly select $k=\text{min}\{b(f)-|M(f)|,|R_{M}(f)|\}$ adjustments from $R_{M}(f)$ for inclusion in $R^{b}_{M}(f)$.
@@
Input: - An instance $I$. - A matching $M$. - Information on the ties in the instance $\phi$.
@@ 5       $R_{M},R^{b}_{M}\mathrel{{:}{=}}\emptyset$;
7             $\xi_{f}\mathrel{{:}{=}}\text{the set of currently unmatched candidates for }f$;
8             for $(each\;x\in\xi_{f})$ do
@@ 4.3 Obtaining a stable matching
Input: - An instance $I$. - A matching $M$. - A tie-breaking strategy $S$. - A set of agents whose preference lists are altered $Q_{a}$. - A base algorithm $A$. - The time threshold for obtaining stable matchings $T$.
@@
3 Function Obtain_Stable_Matching($I,M,S,Q_{a}$):
@@ 6             $y_{worst}\mathrel{{:}{=}}$ the worst partner of $v$ in $M$ under $S$;
[…]
@@
12                   else if (v is full and $(\text{v is full}\textbf{ and }r^{*}_{v}(y)>r^{*}_{v}(y_{worst}))$ then
@@ 26                         $y_{worst}\mathrel{{:}{=}}$ the worst partner of $v$ in $M$ under $S$;
34             $M\mathrel{{:}{=}}$ apply the base algorithm $A$ to find a stable matching after breaking all the ties in the instance $I$ using the tie-breaking strategy $S$;
@@ 4.4 Equity-focused variant of TBLS
By balancing the satisfaction across both sides, these two modifications aim to reduce the sex equality cost while maximizing the matching size. The first modification finds a stable matching with a low sex equality cost after applying the tie-breaking strategy to break all ties. The second modification limits the choice of adjustments to reduce bias toward any one side during the refinement process. We refer to this variant as TBLS-E.
To demonstrate the effectiveness of the second modification, consider the following scenario: suppose the current matching favors the $U$ side, and there is an adjustment $a=(f,x)$, where $f\in U$ and $x\in W$. If $f$ is successfully matched with $x$ following adjustment $a$, then an agent $q\in U$ would be disconnected from $x$. This disconnection may lead $q$ to form a new pairing with another agent $p$, where $r_{q}(p)\geq r_{q}(y)$. Therefore, adjustment $a$ contributes to mitigating the bias toward $U$.
@@ 5 Experiments
In this section, we compare the performance of our TBLS algorithm and its derived version, TBLS-E, with other methods in solving the MAX-SMTI and MAX-HRT problems. The experiments are performed using an Intel Core i9-13900HX CPU (5.40 GHz) with 32GB of RAM. The rest of this section is structured as follows: Section 5.1 introduces the problem set, Section 5.2 details the parameter settings of our algorithms, Section 5.3 presents the performance comparison.
@@ 5.1 Problem set
    For an agent $v$, we begin by generating a random number $10\leq p<$ 1 starting from their first choice.
@@
When we set $g=\text{Geom}(1-p2)$, where $\text{Geom}(1-p2)$ denotes a geometric distribution with the parameter $21-p$ 2, the instances generated by the updated generator are nearly identical to those generated by the original one. In this setting, lower values of $p2$ result in fewer, shorter ties, while higher values lead to a greater number of longer ties. This characteristic simplifies the problem even if $p2$ is high, as nearly all preferences are tied, indicating general indifference. Therefore, we introduce $g=\text{Geom}(p2)$ to generate an inverse relationship between the number of ties and the length of each tie. This modification increases the problem’s complexity when $p2$ is high, as a greater number of shorter ties complicates the solving process.
Similarly, the modified HRT generator accepts five parameters: the number of residents ($n$), the number of hospitals ($m$), the probability of incompleteness ($p1$), the probability of starting a tie ($p2$), and the random generator of the tie length ($g$). Experimental results indicate that HRT problems with capacities uniformly distributed among hospitals present a greater challenge compared to those with randomly allocated capacities [3]. Therefore, the capacity is uniformly distributed among hospitals in the HRT instances.
To evaluate the efficacy of various algorithms in addressing the MAX-SMTI and MAX-HRT problems, we categorize the problems into two sizes: small and large. For the large-sized problems, we set a high value for $p1$ to generate more challenging instances. We systematically vary parameters across both problem sizes, with the configurations detailed in Table 2. In this table, parameters are presented with specific ranges and intervals. For instance, the notation “0–0.9 (0.1)” indicates values ranging from 0 to 0.9 in increments of 0.1. Additionally, the use of commas, such as in “$\text{Geom}(p2),\text{Geom}(1-p2)$”, denotes an alternative parameter set being considered for the parameter. We define each unique combination of $n,m,p1,p2$, $g$ as a configuration. To ensure a robust evaluation, we generate 100 instances for each configuration. The randomly generated instances used in this study are available in the Science Data Bank repository, https://www.scidb.cn/en/s/ZfauMn.
@@
$n$ 100 1000 100 1000
$m$ 10 10–50 (10)
$p1$ 0–0.9 (0.1) 0.95–0.99 (0.01) 0–0.9 (0.1) 0.9
$p2$ 0–1 (0.1) 0–1 (0.1) 0–1 (0.1) 0–1 (0.1)
$g$ $\text{Geom}(p2),$ $\text{Geom}(1-p2)$ $\text{Geom}(p2),$ $\text{Geom}(1-p2)$ $\text{Geom}(p2),$ $\text{Geom}(1-p2)$ $\text{Geom}(p2),$ $\text{Geom}(1-p2)$
@@ 5.2 Parameter settings
In both TBLS and TBLS-E, the parameters $c=0.9$ and $p_{d}=0.05$ are consistently maintained across all configurations. The parameter $c$ is utilized to compute $e_{M}$, which estimates the minimum size of the matching during the search process. Meanwhile, $p_{d}$ denotes the probability of executing a disruption. Furthermore, the time threshold $T$ for obtaining stable matchings is set equal to the time required to obtain a stable matching with the use of the base algorithm at the beginning. The parameters $k_{u}$ and $k_{w}$, which adapt according to the scale and complexity of the problem, are detailed in Table 3. Specifically, $k_{u}$ and $k_{w}$ represent the number of agents selected from sets $U$ and $W$ for disruption, respectively. In the context of the MAX-SMTI problem, $U$ and $W$ represent men and women, respectively, while in the MAX-HRT problem, $U$ and $W$ correspond to residents and hospitals, respectively. For TBLS, GS serves as the base algorithm, whereas for TBLS-E, PDB is used as the base algorithm.
@@ 5.3 Performance comparison
In each configuration, we calculate the average result across 100 instances for each algorithm. An algorithm is deemed the “winner” in a configuration if its result is not worse than that of any other algorithm. To compare the performance of different algorithms, we count the number of wins and compute the overall average result across all configurations. Calculating the overall average result of an algorithm involves two steps. First, the result for each configuration is calculated by averaging the outcomes of 100 instances. Then, the results for each configuration are averaged to obtain the final overall result. These two metrics provide a comprehensive comparison of the overall performance of the algorithms across various configurations. The results for each configuration are available in Online Resource 1.
For the MAX-SMTI problem, we evaluate the solution quality of TBLS and TBLS-E by examining the matching size and the sex equality cost (SECost). These results are compared against those of AS, GSA2, GSM, LTIU, and MCS. Additionally, we compare the execution time of TBLS and TBLS-E with that of AS, LTIU, and MCS, as they are all local search algorithms. For the MAX-HRT problem, we assess the solution quality of TBLS, focusing on the matching size, in comparison with ASBM, HA, HPA, HR, and MCA. Furthermore, we compare the execution time of TBLS with that of HR and MCA, since they all employ local search techniques.

### v2

@@ 8 $iter\mathrel{{:}{=}}0$;
10    $S,Q_{a}\mathrel{{:}{=}}$ RefineTieBreakingStrategy($I,M,S$);
11    $M\mathrel{{:}{=}}$ ObtainStableMatching($I,M,S,Q_{a}$);
@@ 4.1 Evaluation function
where $bigM$ is a large enough number to be fixed later, $|L_{v}|$ represents the length of agent $v$’s preference list, and b(v)|M(v)|b(v)-|M(v)| indicates the number of unassigned positions of agent $v$.
@@
where $e_{M}$ denotes the estimated minimum size of the matching throughout the entire search process. To simplify parameter settings, we define $e_{M}=c\cdot|M_{init}|$, where $c$ represents a ratio varying between 0 and 1, and $M_{init}$ denotes the matching obtained from the random tie-breaking strategy at the beginning.
@@ 4.2 Refining the tie-breaking strategy
A tie-breaking strategy, denoted as $S$, converts the original preference lists $L$, which contain ties, into tie-free (i.e., strictly ordered) preference lists $L^{*}$. Let $r^{*}_{v}(x)$ represent the rank of $x$ in $v$’s tie-free preference list. In the subroutine RefineTieBreakingStrategy, outlined in Algorithm 2, the tie-breaking strategy is updated as follows:
@@
    Obtain a set of candidate adjustments $R^{b}_{M}$ via the subroutine ObtainAdjustments (line 3).
@@
Additionally, a set $Q_{a}$ is maintained to record the agents whose preference lists have been altered during the refinement process. This set will subsequently be used in the BP removal process within the subroutine ObtainStableMatching, which is discussed in Section 4.3.
@@
Input: - An instance II. - A matching $M$. - A tie-breaking strategy $S$. - Information on the ties in the instance $\phi$. - A small probability $p_{d}$. - Two integer $k_{u}$ and $k_{w}$.
@@
3 Function RefineTieBreakingStrategy($I,M,S$):
@@
4 $R^{b}_{M}\mathrel{{:}{=}}$ ObtainAdjustments($I,M$);
[…]
@@ 8       $Q_{a}\mathrel{{:}{=}}Q_{a}\cup v$;
9       for (each tie in v’s preference list)($\text${each tie in v's preference list}) do
10            Break the tietie arbitrarily and update $S$;
@@ 18    $Q_{a}\mathrel{{:}{=}}Q_{a}\cup x$;
Next, we detail the process of identifying adjustments that can increase the likelihood of matching free agents and the construction of $R^{b}_{M}$. In the subroutine ObtainAdjustments, as illustrated in Algorithm 3, $R^{b}_{M}$ is generated through the following two steps:
@@
    Identify a set $R_{M}$ of all potential adjustments (lines 4-11). Given a stable matching $M$, consider each free agent $f\in F_{M}$. For every agent $x\in\xi_{f}^{M}$, if the set intersection $\phi_{f}(x)\cap M(x)$ is non-empty, an adjustment is initiated. This adjustment aims to improve $f$’s rank in $x$’s tie-free preference list by positioning $f$ at the highest rank among those agents in $\phi_{f}(x)$, thereby ensuring $f$ ranks above at least one agent currently matched with $x$. This adjustment introduces $(f,x)$ as a BP and enhances $f$’s likelihood of being matched with $x$. All adjustments in $M$ are stored in $R_{M}$, which can be represented as $R_{M}=\{(f,x)\mid f\in F_{M},x\in\xi_{f}^{M},\phi_{f}(x)\cap M(x)\neq\varnothing\}$. The set of adjustments related to one free agent $f$ is denoted by $R_{M}(f)=\{a\in R_{M}\mid a[1]=f\}$, where $a[i]$ is the $i$-th element in adjustment $a$. Additional discussions on the relationship between our proposed tie-adjustment mechanism and the well-established breakmarriage operation are provided in Section S2 of Online Resource 1.
@@
    Construct $R^{b}_{M}$ by sampling from $R_{M}$ (lines 12-16). One free agent may have multiple adjustments, and the number of adjustments can vary among different free agents. In some cases, agents with many unassigned positions might have relatively few adjustments, while those with fewer unassigned positions might have many adjustments. This imbalance can impact the performance of the algorithm, particularly when it is difficult to increase the overall matching size by filling the positions of those agents with fewer unassigned positions. Consequently, a balanced set of adjustments, $R^{b}_{M}$, is constructed through sampling from $R_{M}$. Specifically, for each free agent $f\in F_{M}$, we randomly select $k=\text{min}\{b(f)-|M(f)|,|R_{M}(f)|\}$ adjustments from $R_{M}(f)$ for inclusion in $R^{b}_{M}(f)$. The value of $k$ is chosen to ensure that the number of candidate adjustments for each free agent $f$ does not exceed its unfilled capacity, which is defined as b(f)|M(f)|b(f)-|M(f)|.
@@
Input: - An instance II. - A matching $M$. - Information on the ties in the instance $\phi$.
@@ 5 $R_{M},R^{b}_{M}\mathrel{{:}{=}}\emptyset$;
7    $\xi_{f}^{M}\mathrel{{:}{=}}\text{the set of currently unmatched candidates for }f$;
8    for $(each\;x\in\xi_{f}^{M})$ do
@@ 4.3 Obtaining a stable matching
Input: - An instance II. - A matching $M$. - A tie-breaking strategy $S$. - A set of agents whose preference lists are altered $Q_{a}$. - A base algorithm AA. - The time threshold for obtaining stable matchings $T$.
@@
3 Function ObtainStableMatching($I,M,S,Q_{a}$):
@@ 6    $y_{worst}\mathrel{{:}{=}}$ the worst partner of $v$ in $M$ under $S$;
8    for $(each\;y\in v^{\prime}s\;\text{tie-free preference list})$ do
@@
[…]
@@ 26          $y_{worst}\mathrel{{:}{=}}$ the worst partner of $v$ in $M$ under $S$;
34    $M\mathrel{{:}{=}}$ apply the base algorithm $A$ to find a stable matching after breaking all the ties in the instance II using the tie-breaking strategy $S$;
@@ 4.4 Equity-focused variant of TBLS
By balancing the satisfaction across both sides, these two modifications aim to reduce the sex equality cost while maximizing the matching size. The first modification finds a stable matching with a low sex equality cost after applying the tie-breaking strategy. The second modification limits the choice of adjustments to reduce bias toward any one side during the refinement process. We refer to this variant as TBLS-E.
To demonstrate the effectiveness of the second modification, consider the following scenario: suppose the current matching favors the $U$ side, and there is an adjustment $a=(f,x)$, where $f\in U$ and $x\in W$. If $f$ is successfully matched with $x$ following adjustment $a$, then an agent $q\in U$ would be disconnected from $x$. This disconnection may lead $q$ to form a new pairing with another agent $p$, where $r_{q}(p)\geq r_{q}(x)$. Therefore, adjustment $a$ contributes to mitigating the bias toward $U$.
@@ 5 Experiments
In this section, we compare the performance of our TBLS algorithm and its derived version, TBLS-E, with other methods in solving the MAX-SMTI and MAX-HRT problems. The experiments are performed using Python on a machine with an Intel Core i9-13900HX CPU (5.40 GHz) and 32GB of RAM. The source code is available at: https://github.com/Junyuan-Qiu/Stable-Matching-Local-Search. The rest of this section is structured as follows: Section 5.1 introduces the problem set, Section 5.2 details the parameter settings of our algorithms, Section 5.3 presents the performance comparison.
@@ 5.1 Problem set
    For an agent $v$, we begin by generating a random number $0\leq p<1$ starting from their first choice.
@@
When we set $g=\text{Geom}(1-p_{2})$, where $\text{Geom}(1-p_{2})$ denotes a geometric distribution with the parameter $1-p_{2}$, the instances generated by the updated generator are nearly identical to those generated by the original one. In this setting, lower values of $p_{2}$ result in fewer, shorter ties, while higher values lead to a greater number of longer ties. This characteristic simplifies the problem even if $p_{2}$ is high, as nearly all preferences are tied, indicating general indifference. Therefore, we introduce $g=\text{Geom}(p_{2})$ to generate an inverse relationship between the number of ties and the length of each tie. This modification increases the problem’s complexity when $p_{2}$ is high, as a greater number of shorter ties complicates the solving process.
Similarly, the modified HRT generator accepts five parameters: the number of residents ($n$), the number of hospitals ($m$), the probability of incompleteness ($p_{1}$), the probability of starting a tie ($p_{2}$), and the random generator of the tie length ($g$). Experimental results indicate that HRT problems with capacities uniformly distributed among hospitals present a greater challenge compared to those with randomly allocated capacities [3]. Therefore, the capacity is uniformly distributed among hospitals in the HRT instances.
To evaluate the efficacy of various algorithms in addressing the MAX-SMTI and MAX-HRT problems, we categorize the problems into two sizes: small and large. For the small-sized problems, we select values of $p_{1}$ that result in instances which are difficult to solve by simply applying the GS algorithm with randomly broken ties. For the large-sized problems, we set a high value for $p_{1}$ to generate more challenging instances. We systematically vary parameters across both problem sizes, with the configurations detailed in Table 2. Furthermore, we conduct scalability experiments to assess how the performance of our algorithms scales as the input size increases from 1,000 to 10,000, with the configurations provided in Table 3. In these two configuration tables, parameters are presented with specific ranges, intervals, or formulas. For instance, the notation “0.7–0.9 (0.1)” indicates values ranging from 0.7 to 0.9 in increments of 0.1. The use of commas, such as in $\text{Geom}(p_{2}),\text{Geom}(1-p_{2})$”, denotes an alternative parameter set being considered for the parameter. Additionally, the formula “(n10)/n(n-10)/n” represents a value that is computed based on $n$. We define each unique combination of $n,m,p_{1},p_{2}$, $g$ as a configuration. To ensure a robust evaluation, we generate 100 instances for each configuration. The randomly generated instances used in this study are available in the Science Data Bank repository, https://www.scidb.cn/en/s/ZfauMn.
@@
$n$ 100 1,000 100 1,000
$m$ - - 10 10–50 (10)
$p_{1}$ 0.7–0.9 (0.1) 0.95–0.99 (0.01) 0.3–0.9 (0.1) 0.9
$p_{2}$ 0.1–1 (0.1) 0.1–1 (0.1) 0.1–1 (0.1) 0.1–1 (0.1)
$g$
$\text{Geom}(p_{2}),$
$\text{Geom}(1-p_{2})$
$\text{Geom}(p_{2}),$
$\text{Geom}(1-p_{2})$
$\text{Geom}(p_{2}),$
$\text{Geom}(1-p_{2})$
$\text{Geom}(p_{2}),$
$\text{Geom}(1-p_{2})$
[…]
Table 3: Problem configurations for the scalability experiments
Parameter MAX-SMTI MAX-HRT
$n$ 1,000-10,000 (1,000) 1,000-10,000 (1,000)
$m$ - 30
$p_{1}$ (n10)/n(n-10)/n 0.9
$p_{2}$ 0.5 0.5
$g$ $\text{Geom}(p_{2})$ $\text{Geom}(p_{2})$
@@ 5.2 Parameter settings
In both TBLS and TBLS-E, the parameters $c=0.9$ and $p_{d}=0.05$ are consistently maintained across all configurations. The parameter $c$ is utilized to compute $e_{M}$, which estimates the minimum size of the matching during the search process. Meanwhile, $p_{d}$ denotes the probability of executing a disruption. Furthermore, the time threshold $T$ for obtaining stable matchings is set equal to the time required to obtain a stable matching with the use of the base algorithm at the beginning. The parameters $k_{u}$ and $k_{w}$, which are adapted according to the scale and complexity of the problem, are detailed in Table 4. Specifically, $k_{u}$ and $k_{w}$ represent the number of agents selected from sets $U$ and $W$ for disruption, respectively. In the context of the MAX-SMTI problem, $U$ and $W$ represent men and women, respectively, while in the MAX-HRT problem, $U$ and $W$ correspond to residents and hospitals, respectively. For TBLS, GS serves as the base algorithm, whereas for TBLS-E, PDB is used as the base algorithm.
@@ 5.3 Performance comparison
In each configuration, we calculate the average result across 100 instances for each algorithm. Specifically, for each algorithm and under each configuration, the average performance metrics reported in this study are computed over 100 instances using the following equations:
$\text{Average number of singles / unassigned positions}=n-\frac{1}{|S|}\sum_{M_{i}\in S}|M_{i}|$ (1)
$\text{Average SECost}=\frac{1}{|S|}\sum_{M_{i}\in S}\text{SECost}(M_{i})$ (2)
$\text{Average execution time}=\frac{1}{|I|}\sum_{i=1}^{|I|}\text{time}(i)$ (3)
where $I$ is the set of instances in the configuration, $M_{i}$ is the resulting matching produced by the algorithm when solving instance $i$, and $S$ denote the set of stable matchings among these, i.e., $S=\{M_{i}\mid M_{i}\text{ is stable for }i=1,2,\cdots,|I|\}$.
Note that all local search algorithms used to compare the quality of the solution produced stable matches on all small-sized and large-sized instances, except that MCA solved only 99 of 100 instances in one configuration. For this reason, we did not report the percentage of stable matchings in the results.
An algorithm is deemed the “winner” in a configuration if its result is not worse than that of any other algorithm. To compare the performance of different algorithms, we count the number of wins and compute the overall average result across all configurations. Calculating the overall average result of an algorithm involves two steps. First, the result for each configuration is calculated by averaging the outcomes of 100 instances. Then, the results for each configuration are averaged to obtain the final overall result. These two metrics—the number of wins and the overall average—offer a comprehensive basis for comparing the performance of the algorithms across different configurations.
For the MAX-SMTI problem, we evaluate the solution quality of TBLS and TBLS-E by examining the matching size and the sex equality cost (SECost). These results are compared against those of AS, GSA2, GSM, LTIU, MCS, and PDB. Specifically, LTIU is built upon the SML2 algorithm detailed in [12], and PDB is applied to obtain a stable matching after randomly breaking all the ties. Additionally, we compare the execution time of TBLS and TBLS-E with that of AS, LTIU, and MCS, as they are all heuristic search methods. For the MAX-HRT problem, we assess the solution quality of TBLS, focusing on the matching size, in comparison with ASBM, HA, HPA, HR, and MCA. Furthermore, we compare the execution time of TBLS with that of HR and MCA, since they all employ heuristic techniques. For both problems, we do not include approximation algorithms such as GSA2, HA, etc., in the execution time comparison, since they operate under fundamentally different assumptions and performance objectives compared with heuristic algorithms. The detailed results for each configuration presented in Sections 5.3.1 and 5.3.2 are provided in Online Resource 2, whereas the results corresponding to Section 5.3.3 are included in Online Resource 3.